OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16 sq. in

OpenStudy (anonymous):

16=4pir^2

OpenStudy (anonymous):

4=pi(r^2)

OpenStudy (anonymous):

i ended up with 64/3pi using the same method outkast seems to be

OpenStudy (anonymous):

4/3 π r3= volume 4*pi*r^2= surface area r=2 answer: 32/3 PI

OpenStudy (anonymous):

thats the only way i can think of unless there is some sort of volume to surface area ratio

OpenStudy (anonymous):

r = sqrt(4/pi)

OpenStudy (anonymous):

surface area is 16 square inches not 16pi... if hombre did what i was saying then i believe that would be reight

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

64/3 * sqrt(4/pi) silly algebra mistake

OpenStudy (anonymous):

\[\sqrt{4/\pi}=]=2\sqrt{\pi}/\pi\]

OpenStudy (anonymous):

thanks